Value-Stream Costing Objective

During the week of June 12, Harrison Manufacturing produced and shipped 18,200 units of its aluminum wheels: 4,400 units of Model A and 13,800 units of Model B. The following costs were incurred:

Materials        Salaries/
Wages
       Machining        Other        Total Cost
Order processing        $18,800                      $18,800
Production planning        136,800                      136,800
Purchasing        26,100                      26,100
Stamping $355,000        35,500        $37,600        $19,600        447,700
Welding 160,000        40,000        40,000        12,000        252,000
Cladding 390,000                             390,000
Testing        11,000                      11,000
Packaging and shipping        9,000                      9,000
Invoicing        12,400                      12,400
  Total $905,000        $289,600        $77,600        $31,600        $1,303,800

Required:

1. Assume initially that the value-stream costs and total units shipped apply only to one model (a single-product value stream). Calculate the unit cost. Round your answer to the nearest dollar amount.
$ per unit

2. Calculate the unit cost for Models A and B. Round your answers to the nearest dollar amount.

Unit Cost
Model A $
Model B $

3. What if Model A is responsible for 40 percent of the materials cost? Calculate the unit materials cost for Models A and B. Round your answers to the nearest dollar amount.

Unit Cost
Model A $
Model B $

Calculate the total unit cost for Models A and B. Round your interim calculations and final answers to the nearest dollar amount.

Unit Cost
Model A $
Model B $
Answer #1
Part- A & B
Unit Cost := Total Cost / Total Unit
=1303800/(4400+13800)= $71.64
Eac Model A & B will be the same cost $71.64
Part-C
Model A ( Unit Material Cost )= Total Cost/ Total Unit  
(905000*40%)/4400= $82.27
Model B ( Unit Material Cost )= Total Cost/ Total Unit  
(905000*60%)/13800= $39.35
Computation of Connversion Cost
Unit Conversion Cost =Total Cost/ Total Unit
=(1303800-905000)/(4400+13800)= $21.91
Computation of Total Unit cost
Particular Material Cost Conversion Cost Total
Model A $82 $22 $104
Model B $39 $22 $61